Fermi National Laboratory


Accelerator Update for January 14, 2005 - January 17, 2005

Friday January 14
The day shift began with Operations monitoring store 3920. The Main Injector (MI) was off due to the MI-60 power supply/VCB problem. Operations monitored the Antiprotons Source (Pbar) stack. MiniBooNE and the Meson experiments MCenter and MTest were also off.

Operations established beam to MiniBooNE at 8:12 AM.

At 3:14 PM, a Linac expert reported finding a broken BPM that was used for the 400 MeV area tuning program. The program can't be used until the BPM gets repaired.

Operations reported at 5:17 PM that they received permission to operate MI.

Operations established an 8 GeV DC ramp in MI at 5:45 PM.

Operations established 120 GeV beam to the MI dump at 7:15 PM.

Pbar began stacking at 7:22 PM.

Operations established beam to the Meson dump at 8:11 PM.

Operations established reverse protons to the Recycler at 8:21.

Operations terminated store 3920 at 8:52 PM.

Operations put the Tevatron (TeV) through a dry squeeze at 9:35 PM and then into a wet squeeze.

Saturday January 15
The midnight shift began with Pbar stacking and beam to MiniBooNE.

Operations put the TeV into shot setup at 12:28 AM.

At 2:40 AM, a Pbar RF station (ARF1) tripped off while unstacking antiprotons for transfer to the TeV. About 23mA of antiprotons were lost. Later, operators discovered that the real problem was elsewhere.

Operations established store 3923 at 3:16 AM with an initial luminosity of 44.74E30. Pbar resumed stacking.

At 4:18 AM, a TeV separator (A49H) spark counter alarmed, but it didn't seem to cause a luminosity problem.

Operations resumed sending beam to MCenter at 5:20 AM.

Operations put the Recycler into shot setup at 7:42 AM. By 9:37 AM, the Recycler had stashed six shots of antiprotons. Pbar resumed stacking at 9:52 AM.

Operations resumed sending beam to MCenter and MTest at 10:21 AM.

Operations terminated store 3923 at 11:03 PM and then put the TeV into shot setup at 11:15 PM.

Sunday January 16
The midnight shift began with the TeV in shot setup and with MiniBooNE taking beam. (Switchyard beam is always turned off during shot setups.)

Operations established store 3924 at 1:38 AM with an initial luminosity of 77.22E30. Pbar resumed stacking.

Operations reestablished beam to MCenter at 2:10 AM.

Operations put the Recycler into shot setup at 8:28 AM and completed the transfers at 10:32 AM. Pbar resumed stacking and MCenter resumed taking beam.

At 12:01 PM, Operations reported a MiniBooNE sump water level alarm.

At 3:52 PM, Operations disabled the MiniBooNE beam due to the water level problem.

Operations put the Recycler into shot setup at 4:16 PM and completed the transfers at 5:21 PM. Pbar resumed stacking and MCenter resumed taking beam.

MiniBooNE began taking beam again. Experts and the Duty Mechanic believed the alarm was due to a frozen discharge pipe.

MTest experimenters reported at 8:51 PM that they were finished with the beam for the night.

At 9:11 PM, the TeV separator A49 tripped off due to a LCW flow monitor. Operators back flushed the monitor and this cleared the alarm, but the TeV luminosity had already taken a hit.

Monday January 17
The Midnight shift began with Operations monitoring store 3924, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:25 AM, a Booster RF station (BRF16) tripped off. It reset okay, but then BRF18 tripped off for the first of three times in about three and a half hours.

Operations terminated store 3924 around 8 AM.

The Plan for Monday
The plan for today is to establish a store and then continue with normal operations.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/18/2005   email Fermilab